# Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices

### v1

@@ III-D2 Parameter Calibration and On-Chip Optimization
where $S_{c}^{(i)}$ is the total spike count for class $c$ on sample $i$, $y_{i}$ is the true label, and $\mathbb{I}[\cdot]$ is the indicator function. This allows data-driven threshold selection for each class. If performance deviation is observed relative to the original CNN, light on-chip fine-tuning is conducted using Akida’s incremental learning capability, adjusting only the last output layer to optimize for domain shift or quantization artifacts.
@@ IV-A Datasets
We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from Hospital Sultanah Bahiyah, Malaysia. Both sets reflect a range of lesion types and clinical diversity.
@@ IV-A1 HAM10000 Dataset
IV-A2 Hospital Sultanah Bahiyah Clinical Dataset
A proprietary clinical dataset was established in partnership with Hospital Sultanah Bahiyah, comprising 3,162 dermatoscopic images from 1,235 patients collected between June 2022 and February 2024. To ensure comparability with benchmark studies, lesion categories were selected to match the seven classes in HAM10000. Rare and unclassified lesions were excluded to maintain consistency in diagnostic labeling and facilitate joint evaluation. Each case was independently annotated by at least two board-certified dermatologists and histopathologically confirmed where possible. All images and metadata were anonymized following institutional ethical guidelines (approved protocol: NJG-2022-3233-CN). Dataset partitioning followed the same 70%/10%/20% train/validation/test split as HAM10000, stratified by disease category. This dataset provides a clinically diverse, real-world validation source for model generalization and robustness.
@@ IV-B Experimental Setup
All model training, validation, and CNN-to-SNN conversion were conducted on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and NVIDIA RTX 3090 GPU, running Ubuntu 22.04 LTS. The software environment included Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1. Neuromorphic inference was performed on a BrainChip Akida AKD1000 PCIe board installed in the same system, with deployment and testing managed via the Akida Python API and default board settings.
@@ IV-C Analysis of Classification Results on HAM10000 Dataset
Table I reports the precision, recall, and accuracy of the proposed model for each class on the HAM10000 test set. The model achieves consistent performance across all lesion categories, including minority classes. Notably, the architecture maintains stable results under limited training data and on-chip incremental learning, enabling effective clinical adaptation in evolving or low-resource scenarios. With a Top-1 accuracy of 91.6% and macro F1 of 82.4%, the system demonstrates effective discrimination of both common and rare lesions.
TABLE I: Class-wise precision, recall, F1 score, and accuracy of QANA on the HAM10000 test set.
Class Precision Recall F1 Accuracy
Actinic keratoses 0.890 0.933 0.911 0.933
Basal cell carcinoma 0.890 0.901 0.896 0.901
Benign keratosis-like lesions 0.866 0.853 0.859 0.853
Dermatofibroma 0.925 0.976 0.950 0.976
Melanocytic nevi 0.887 0.817 0.851 0.817
Vascular lesions 0.949 0.966 0.957 0.966
Melanoma 0.956 0.933 0.944 0.933
Average 0.909 0.911 0.910 0.910
@@ IV-E Inference Speed and Energy Consumption
We quantitatively evaluated the inference latency and energy consumption of our neuromorphic model in comparison with both conventional CNN baselines and other SNN-converted architectures. All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA-RTX-3090-GPU and an Intel Xeon CPU. For each model, the reported inference latency corresponds to the average per-image processing time over 10,000 test samples. Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida. Table III presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations. Our neuromorphic model achieves the lowest inference latency and energy consumption of all evaluated architectures. When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 99.0% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
@@ IV-F Ablation Study of Model Components
Table IV reports the ablation results on the HAM10000 test set, illustrating the contribution of each core module to the overall performance. Modules were incrementally enabled to measure their isolated and cumulative effects on classification metrics, including accuracy, recall, precision, F1 score, and AUC-ROC (%). The results demonstrate that each module provides consistent improvements, with the complete model achieving the highest accuracy and F1 score.
@@ V Conclusion
In this paper, we proposed QANA, a quantization-aware neuromorphic framework for skin lesion classification on edge devices. Extensive experiments on the large-scale HAM10000 benchmark and a real-world clinical dataset show that QANA achieves state-of-the-art accuracy (91.6% Top-1, 82.4% macro F1 on HAM10000; 90.8%/81.7% on the clinical set) while enabling real-time and energy-efficient inference on the BrainChip Akida platform (1.5 ms latency, 1.7 mJ per image). These results demonstrate that QANA is highly effective for portable medical analysis and AI deployment in dermatology under limited computing resources.
@@ VI Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62172336 and 62032018). The authors gratefully acknowledge BrainChip Holdings Ltd. for providing technical support and the Akida AKD1000 hardware platform, whose powerful neuromorphic computing capabilities enabled strong performance of the SNN model. The authors also extend their appreciation to Dr. Atif Mansoor, Dr. Bo Miao and their teams for their preliminary contributions to this research.

### v2

@@ III-D2 Parameter Calibration and On-Chip Optimization
where $S_{c}^{(i)}$ is the total spike count for class $c$ on sample $i$, $y_{i}$ is the true label, and $\mathbb{I}[\cdot]$ is the indicator function. This allows data-driven threshold selection for each class. If performance deviation is observed relative to the original CNN, light on-chip fine-tuning is conducted using Akida’s incremental learning capability, adjusting only the last output layer to optimize for domain shift or quantization artifacts. For higher-detail cases, a $2\times 2$ tiling fallback is triggered when $\max_{c}\hat{p}_{c}<0.60$. Four overlapping $64\times 64$ patches (20% stride overlap) are evaluated and logits averaged. On HAM10000, the trigger rate is 18%.
@@ IV-A Datasets
We used two datasets: the public HAM10000 benchmark and a proprietary clinical dataset from local hospital. Both sets reflect a range of lesion types and clinical diversity.
@@ IV-A1 HAM10000 Dataset
IV-A2 local Clinical Dataset
A non-public clinical dermatoscopic dataset was provided by a partner hospital under a data-use agreement. The dataset comprises 3,162 images from 1,835 patients acquired during routine care between June 2022 and February 2024. No human experiments or prospective data collection were conducted by the authors. Lesion categories were harmonized to the seven HAM10000 classes; rare or unclassified cases were excluded to maintain consistent labeling. Each case was independently annotated by at least two board-certified dermatologists, with histopathology confirmation when available. All records were de-identified before access and used solely for secondary analysis. Dataset partitioning followed the same $70\%/10\%/20\%$ train/validation/test split as HAM10000, stratified by disease category, providing a clinically diverse, real-world validation source for model generalization.
@@ IV-B Experimental Setup
All training and conversion are performed on a workstation with an Intel Core i9-12900K CPU, 128 GB RAM, and an NVIDIA RTX 3090, running Ubuntu 22.04 LTS. Software includes Python 3.9, CUDA 11.8, TensorFlow 2.10, and Akida MetaTF SDK v2.2.1. Unless otherwise stated, batch size is 1 for latency and energy measurement. We use AdamW (weight decay $10^{-4}$), initial learning rate $3\times 10^{-4}$ with cosine decay for 120 epochs, and early stopping on validation macro F1. Quantization-aware training uses 8-bit affine per-tensor quantization for weights and activations with ReLU6 clipping. Neuromorphic inference uses a BrainChip Akida AKD1000 PCIe board with default runtime settings. GPU energy is measured as the board power above idle (averaged by NVML at 20 Hz) multiplied by the per-image inference time; Akida energy is measured analogously by on-board power monitoring. Reported Akida 1.5 ms corresponds to on-chip execution time; host preprocessing and I/O add <0.2<0.2 ms on average and are reported separately when stated.
TABLE I: Class-wise precision, recall, F1, and per-class accuracy on the HAM10000 test set. The macro F1 equals the unweighted mean of per-class F1.
Class Precision Recall F1 Accuracy
Actinic keratoses 0.840 0.820 0.830 0.820
Basal cell carcinoma 0.830 0.820 0.820 0.820
Benign keratosis-like 0.800 0.780 0.790 0.780
Dermatofibroma 0.900 0.860 0.880 0.860
Melanocytic nevi 0.800 0.740 0.770 0.740
Vascular lesions 0.870 0.850 0.860 0.850
Melanoma 0.830 0.810 0.820 0.810
Macro average 0.853 0.826 0.824 0.826
@@ IV-C Analysis of Classification Results on HAM10000 Dataset
Table I reports per-class precision, recall, F1, and per-class accuracy under the $64\times 64$ single-pass setting. The macro F1 is $0.824$. Vascular lesions ($\mathrm{F1}=0.860$) and dermatofibroma ($\mathrm{F1}=0.880$) rank highest, whereas benign keratosis-like lesions ($\mathrm{F1}=0.790$) and melanocytic nevi ($\mathrm{F1}=0.770$) are lower, consistent with known visual overlap between these categories and class imbalance in the dataset. Errors are concentrated between nevi and benign keratosis-like lesions and at the nevi–melanoma boundary. The channel reweighting modules (ECA/SE) suppress redundant responses and improve minority coverage, which aligns with the relatively stronger performance in vascular lesions and dermatofibroma.
Per-class accuracy in the table largely mirrors recall by construction and is provided for completeness. All decisions use plain argmax without threshold tuning. The image-level Top-1 accuracy on the same test split is $91.6\%$, while the macro F1 remains more conservative due to equal weighting across classes.
@@ IV-E Inference Speed and Energy Consumption
We quantitatively evaluated the inference latency and energy consumption of our neuromorphic model in comparison with both conventional CNN baselines and other SNN-converted architectures. All SNN models were deployed on the Akida AKD1000 PCIe board, while CNN baselines were tested on both NVIDIA RTX 3090 GPU and an Intel Xeon CPU. For each model, the reported inference latency corresponds to the average per-image processing time over 1000 test samples. Energy consumption per image was measured as the mean device power during inference multiplied by the average inference time, using on-board power monitoring for Akida and NVIDIA-smi for GPU and Intel RAPL for CPU.
To ensure a fair and hardware-consistent comparison, all tested CNN architectures were converted to SNNs using our pipeline before deployment on Akida. Table III presents a comprehensive summary of inference latency and energy consumption across all evaluated models and platforms, as well as the relative reduction of these metrics for SNNs on Akida compared to their CNN GPU and CPU implementations. Our neuromorphic model achieves the lowest inference latency and energy consumption of all evaluated architectures. When deployed as an SNN on the Akida platform, it completes classification in 1.5 ms per image and requires just 1.7 mJ, representing an 94.6% reduction in latency and over 98.6% reduction in energy compared to the equivalent CNN on GPU surpassing all other state-of-the-art CNN-to-SNN conversion baselines.
For images requiring higher spatial detail, the $\times2$ tiling fallback increases latency and energy approximately in proportion to the number of patches while remaining within the power envelope of the edge platform. This fallback is applied conditionally based on a confidence threshold on the single-pass prediction and does not change the training protocol.
@@ IV-F Ablation Study of Model Components
Table IV summarizes a stepwise ablation in which we start from a compact baseline and incrementally enable Ghost, ECA, SE, quantization-aware head with the same augmentation policy, embedding-space SMOTE, and on-chip incremental learning. All experiments use the same input resolution, optimizer, schedules, and 8-bit quantization configuration to isolate the effect of each component. Metrics are computed on the same test split.
The drop observed for “Baseline + Ghost” is expected under a strictly width-preserving replacement of standard convolutions by Ghost modules. Ghost generates a fraction of channels through inexpensive linear transforms of a reduced set of primary channels. When used alone and early in the stack, this reduces the number of learned base filters that directly capture fine dermatoscopic structures such as pigment networks, streaks, and globules. The cheap transforms introduce correlated feature maps that do not add discriminative power without a mechanism to suppress redundancy. Under 8-bit activation clipping, these correlated responses compress the effective dynamic range and make the batch-norm folding more sensitive to local contrast, which in turn degrades minority-class recall. The combination of early downsampling and reduced primary channels further limits high-frequency detail preserved from $64\times 64$ inputs. As a result, the test accuracy decreases relative to the baseline when Ghost is introduced in isolation.
Adding ECA recovers and surpasses the baseline by providing lightweight channel recalibration that suppresses redundant ghosted maps and reweights informative channels. The subsequent SE block contributes a complementary global channel prior that stabilizes inter-class boundaries and improves calibration. Enabling the quantization-aware head with the same augmentation policy improves the match between integer ranges at the readout and the upstream bounded activations, which tightens the loss landscape after BN folding. Injecting SMOTE in the embedding space raises coverage of minority manifolds without perturbing low-level image statistics and yields consistent gains in macro metrics. Finally, on-chip incremental learning adapts the last-stage parameters to post-conversion activation statistics and to the deployment data distribution, producing the highest accuracy and F1. We also observed that widening the primary branch of Ghost or delaying its use to later stages can mitigate the initial drop, but we kept the compact configuration to satisfy on-chip memory and latency constraints and relied on attention and calibration to recover accuracy.
@@ V Conclusion
We presented QANA, a quantization-aware neuromorphic pipeline for skin lesion classification on edge hardware. Under a fixed $64\times 64$ input and identical preprocessing across methods, QANA reaches 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical dataset. The compiled SNN runs at 1.5 ms on chip and 1.7 mJ per image on Akida, measured as power above idle. These results indicate that conversion-preserving design and deployment-aware training enable accurate and efficient on-device dermatology inference.
@@ VI Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 62172336 and 62032018). The authors gratefully acknowledge BrainChip Holdings Ltd. for providing technical support and the Akida AKD1000 hardware platform, whose powerful neuromorphic computing capabilities enabled strong performance of the SNN model. The authors also extend their appreciation to Dr. Atif Mansoor and his teams for their preliminary contributions to this research.